# On the comparison of ultradifferentiable weight function settings

### v1

@@ Abstract.
We summarize, revisit and compare different notions of ultradifferentiable settings defined in terms of weight functions. More precisely, we emphasize the difference between the original Beurling-Björck-setting, when measuring the growth of the Fourier-transform, and the more recent approach by Braun, Meise and Taylor, when controlling the growth of the derivatives similarly like in the weight sequence framework. For the equality of both settings the convexity of the defining weight function is required and, somehow, has become a standard assumption. We study the meaning of this convexity condition and its failure in detail. First, we characterize the inclusion relations and hence the equality for Beurling-Björck-type classes in terms of weight functions analogously as it has already been done for the Braun-Meise-Taylor setting and work within the general weight matrix setting. Finally, we construct a technical (counter-)example which is a weight in the sense of Beurling-Björck but violates the convexity condition and which illustrates the different flavor of both weight function settings.
@@ 1. Introduction
When dealing with ultradifferentiable function classes and analogously defined weighted spaces, in the literature “classically” one can find two settings: either using a weight sequence $\mathbf{M}$ or a weight function $\omega$. The ultradifferentiable weight sequence case is more original, we refer to [15] and [16]. On the other hand, in the weight function setting first the growth of the Fourier transform (of compactly supported functions) has been measured; see [1], [3], [23] and Section 6.3 for a summary. Here the expression measuring the admissible growth is given by $\omega$ directly (more precisely by $\exp\circ\omega$) and, indeed, non-trivial weighted test function spaces are treated (“Beurling-Björck setting”). However, in these original frameworks slightly different notions for being a weight function have been considered; we refer to Sections 2.8 and 6.3.
@@
Later, in [7] the nowadays frequently used definition involving the Legendre-Fenchel-Young conjugate $\varphi^{*}_{\omega}$ of $\varphi_{\omega}:t\mapsto\omega(e^{t})$ has been given (“Braun-Meise-Taylor setting”). For such spaces the derivatives of the functions under investigation are weighted analogously to the weight sequence framework; see [7, Lemma 3.3, Prop. 3.4].
@@
In each approach, basic growth and regularity assumptions on the weights $\omega$ are required and one can distinguish between two types: The Roumieu-type, involving the existence quantifier, and the Beurling-type when using the universal quantifier. As it has been mentioned in the introduction in [7], when comparing the assumptions between both settings it follows that requirement subadditivity for $\omega$ in [1], [3] and [23] can be replaced resp. relaxed by $\omega(2t)=O(\omega(t))$. This condition is frequently abbreviated by $(\alpha)$ and we write $(\omega_{1})$ for it; this abbreviation is motivated by the notation used in [29]. A further achievement is that the definition of weighted spaces within the Braun-Meise-Taylor setting admits the possibility to introduce and investigate the notion of quasianalyticity as well analogously to the weight sequence setting (“Denjoy-Carleman-Theorem”).
@@
On the other hand, for establishing this equivalent description in [7, Lemma 3.3, Prop. 3.4] crucially it has been used that $\varphi_{\omega}$ is convex; usually abbreviated by $(\delta)$ and we write $(\omega_{4})$. This has become a standard assumption on $\omega$ also in related weighted settings since then; e.g. when considering globally defined functions of Gelfand-Shilov-type. $(\omega_{4})$ is crucial for obtaining the coincidence of both settings since the proof of the crucial Paley-Wiener-type theorem involves the double conjugate $\varphi^{**}_{\omega}$ and the identity $\varphi^{**}_{\omega}=\varphi_{\omega}$ which holds if and only if $\varphi_{\omega}$ is convex. In general, one only has $\varphi^{**}_{\omega}\leq\varphi_{\omega}$ and for this note that $\varphi^{**}_{\omega}$ and $\varphi^{*}_{\omega}$ both are automatically convex; see the comments in [29, Sect. 3.1] and more details are given in Section 2.9.
@@
In [29] and [24] we have shown that ultradifferentiable classes in the Braun-Meise-Taylor setting can equivalently be described by using the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j))$, $j\in\mathbb{N}$; we refer to the main statement [24, Thm. 5.14] and for this alternative description $(\omega_{1})$ is crucial. It is immediate by definition that properties for $\omega$ can be transferred to $\mathcal{M}_{\omega}$ but, however, for the converse and hence when establishing a characterization of growth conditions, then again $(\omega_{4})$ is becoming relevant in order to treat the double conjugate. Another crucial property for which $(\omega_{4})$ is unavoidable is the equivalence between $\omega$ and each associated weight function $\omega_{\mathbf{W}^{(\ell)}}$; see [24, Lemma 5.7], [29, Theorem 4.0.3, Lemma 5.1.3], and also [13, Lemma 2.5]. On the other hand, for the sake of completeness we mention that basically $\mathcal{M}_{\omega}$ can be introduced for more general $\omega$. Indeed, it suffices to assume that $\omega$ is non-decreasing and $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$ (abbreviated by $(\omega_{3})$ in the following) in order to ensure well-definedness of the Young conjugate and hence of the sequences $\mathbf{W}^{(\ell)}$; we refer to the comments in [29, Sect. 3.1], [26, Sect. 2.7, (a)(c)(a)-(c)], [28, Sect. 6 $(III)$] and the summary in Section 2.9. Via $(\omega_{1})$ also the equality between the weighted spaces holds when the classes are introduced directly (by definition) via weighting the derivatives in terms of $\varphi^{*}_{\omega}$. Indeed, for this equality estimate [24, $(5.10)$] is required and for its proof $(\omega_{4})$ is not needed. For future reference we call any non-decreasing function satisfying $(\omega_{3})$ matrix admissible.
The main aim of this article is to illustrate the difference between the Beurling-Björck and the Braun-Meise-Taylor setting in view of the convexity for $\varphi_{\omega}$. Without this assumption the main Paley-Wiener-type result from [7] and hence the equality between both settings is becoming unclear. Therefore, when treating the pure Beurling-Björck framework the specific role resp. the loss of $(\omega_{4})$ has to be investigated. This question has been motivated by the explanations above and it is based on the comments in [29, Sect. 3.1]. But also in the recent article [20] in the introduction this convexity condition, its possible failure and its role in proofs within the Beurling-Björck are discussed. ($(\omega_{4})$ is also denoted by $(\delta)$ in [20].)
More precisely, by taking into account the fact that $(\omega_{4})$ is in general not preserved under equivalence of weight functions, we can see the first main question:
@@
    How restrictive is $(\omega_{4})$? Does for any weight function (in the sense of Beurling-Björck) an equivalent function exist which satisfies $(\omega_{4})$?
Concerning the notion of equivalence of weight functions and more relevant growth relations we refer to Section 2.3. Connected to $(Q1)$ is the following problem:
@@
    Characterize the inclusion relation, and hence the equality, of the spaces in the Beurling-Björck setting (for both types) in terms of the defining weights. One should expect the same growth relation between weight functions as in the Braun-Meise-Taylor setting (expressed by an O-growth restriction, written $\preceq$) and thus equivalence of weight functions characterizes the equality (as l.c.v.s).
@@
We treat and answer both questions in this work in detail: The last Section 7 is dedicated to $(Q1)$ and to the construction of a (counter-)example showing that there exist weights in the sense of Beurling-Björck such that none equivalent (weight) function satisfies $(\omega_{4})$. Indeed, we even verify that there exist “many” of such weight functions; i.e. uncountable infinitely many different equivalence classes. The geometric construction also involves recent knowledge about the growth index $\gamma(\omega)$ for weight functions $\omega$ introduced in [12]; see Section 2.7. An immediate consequence of this example is to see the limitation of the weight sequence setting in this approach; i.e. when considering associated weight functions $\omega_{\mathbf{M}}$ with $\mathbf{M}$ being a sequence satisfying basic growth and regularity properties. This is due to the fact that $\omega_{\mathbf{M}}$ satisfies automatically $(\omega_{4})$; see Remark 7.6 for more details. On the other hand $\omega_{\mathbf{M}}$ appears frequently in the Braun-Meise-Taylor setting as a (counter-)example and in the comparison results from [6]; this fact again illustrates the difference between both settings. Finally, the constructed $\omega$ in Section 7 is matrix admissible.
$(Q2)$ is studied in detail in Sections 3, 4, 5 and 6; for the main characterizing results in this context see Theorem 4.1 for the Beurling-type and Theorem 4.2 for the Roumieu-type. We treat a very general setting: First, we work with weight function matrices $\mathcal{W}:=\{\omega^{\ell}:\ell>0\}$ with matrix parameter $\ell$, see Definition 3.1. Second, in Sections 3, 4 we are concerned with weighted subspaces of $L^{p}(\mathbb{R}^{d})$, $1\leq p\leq\infty$, subject to $\mathcal{W}$. Indeed, the weights are given by $x\mapsto\exp(\omega^{\ell}(x))$ and so involving the exponential function. Section 5 summarizes the situation when we focus on two special but natural weight function matrices for which the crucial information can be expressed in terms of a single weight function $\omega$: matrices of “exponential-type” and of “dilatation-type”. In Section 6 we apply the information to related and analogous situations: First, in Section 6.1 we consider the setting when weighting a fixed subspace $\mathcal{A}^{p}\subseteq L^{p}(\mathbb{R}^{d})$. In Section 6.2 we focus in detail on the case when weighting the “Fourier image” of a given (fixed) function space $\mathcal{A}$; i.e. not the growth of $f$ directly is controlled but of its Fourier transform $\widehat{f}$. Of course, for this case it is crucial to ensure that the Fourier transform is well-defined on $\mathcal{A}$ and its image is non-trivial. Finally, in each instance we also treat the anisotropic setting; i.e. when the weight functions are not radially extended to $\mathbb{R}^{d}$.
The known Beurling-Björck setting corresponds now to the special weight matrix $\omega^{\ell}:=\ell\omega$, $\ell>0$, where $\omega$ is given and fixed (“exponential type”), see Section 5.1, and when weighting the Fourier image of fixed subspaces (of functions having compact support) in $L^{1}(\mathbb{R}^{d})$. We refer to Section 6.3 for more details.
@@
Before starting with this approach, in the first Section 2 we list, summarize and compare in detail the numerous growth conditions and relations on weight functions in the different settings. One can expect that this content is also interesting for other different weighted settings since we exclusively deal with growth and regularity properties for functions. The same comment applies to the construction provided in Section 7.
@@ 2.2. Relevant growth properties
The next condition is crucial for characterizing desired stability properties for Braun-Meise-Taylor classes (of both types), see [24], [25] and the citations therein:
@@
In [29, $(3.8.4)$] this requirement has been denoted by $(\omega_{1^{{}^{\prime}}})$ but frequently in the literature it is abbreviated by $(\alpha_{0})$. Finally, let us introduce the following two requirements appearing in the Beurling-Björck framework (see [3]):
Finally, the Legendre-Fenchel-Young conjugate (of $\varphi_{\omega}$) is defined as follows:
This transform is crucial when defining weighted classes in the sense of Braun-Meise-Taylor and is also basic for introducing the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}=(W^{(\ell)}_{j})_{j\in\mathbb{N}}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp\left(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j)\right)$. We investigate this transform in more detail in Section 2.9.

### v2

# On inclusion relations of weighted $L^{p}$-type spaces defined in terms of weight function matrices

@@ Abstract.
We introduce new weighted $L^{p}$-type spaces defined in terms of weight function matrices and characterize the inclusion relations in terms of the defining matrices. Moreover, we provide a detailed study concerning the coincidence with the common (non-weighted) $L^{p}$-spaces, the (non-)triviality of such weighted spaces and investigate their translation invariance. The obtained results are then applied to particular weight function matrices which are expressed in terms of one single weight function and a positive real parameter. Also variations of this new weighted setting are discussed; more precisely weighted Banach (sub-)spaces of $L^{p}$ and when weighting the Fourier image of appropriate Banach spaces of functions. The general framework allows to describe the known ultradifferentiable weight function setting by Beurling-Björck which is more original than the approach presented by Braun, Meise and Taylor. When applying the characterization of the inclusion relations to Beurling-Björck-type spaces we are able to emphasize the difference between both ultradifferentiable weight function settings: We construct a technical (counter-)example which is a weight in the sense of Beurling-Björck but violates the standard and crucial convexity condition needed in the Braun-Meise-Taylor setting.
@@ 1. Introduction
When dealing with ultradifferentiable function classes and analogously defined weighted spaces, in the literature “classically” one can find two settings: either using a weight sequence $\mathbf{M}$ or a weight function $\omega$. The ultradifferentiable weight sequence case is more original, we refer to [15] and [16]. On the other hand, in the weight function setting there exist two methods: First the growth of the Fourier transform $\widehat{f}$ (of compactly supported functions $f$) has been measured w.r.t. weighted $L^{1}$ or $L^{\infty}$-based norms; see [1], [3], [23], the explanations in [7] and Section 6.3 for precise definitions and a summary. More precisely, the boundedness of $\int_{\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}dx$ or $\sup_{x\in\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}$ for some $\ell>0$ (“Roumieu-type”) resp. for all $\ell>0$ (“Beurling-type”) is required and so the admissible growth is controlled by $\omega$ directly (“Beurling-Björck setting”).
@@
Later, in [7] the nowadays frequently used definition has been given: Let $\varphi_{\omega}:t\mapsto\omega(e^{t})$ and let $\varphi^{*}_{\omega}$ be the Legendre-Fenchel-Young conjugate given by
Then the derivatives of the functions under investigation are weighted analogously to the weight sequence framework and the weights are given by $W^{(\ell)}_{j}:=\exp(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j))$, $j\in\mathbb{N}$ (“Braun-Meise-Taylor setting”). Here, $\ell>0$ denotes the crucial parameter, hence analogously one has the Roumieu-type $\mathcal{E}_{\{\omega\}}$ and the Beurling-type $\mathcal{E}_{(\omega)}$. For obtaining the coincidence of both weight function settings a crucial Paley-Wiener-type theorem has been shown, see [7, Lemma 3.3, Prop. 3.4].
In each approach basic growth and regularity assumptions on the weights $\omega$ are required and, as it has been mentioned in the introduction in [7], several crucial differences between these two settings can be recognized. For the convenience of the reader we recall now some definitions; see also Sections 2.2, 2.8 and 6.3 for more details and note that the abbreviations of the conditions have already been used in [28].
Definition 1.1.
    continuous,
    non-decreasing,
    $\omega(0)=0$, and
    $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$.
Note: For some considerations an even more general setting is sufficient; occasionally in [12] we have even worked without the assumptions continuity and $\omega(0)=0$.
Definition 1.2.
  • $(i)$
    continuous,
  • $(ii)$
    $\omega(0)=0$,
  • $(iii)$
    $\omega$ satisfies $(\omega_{\operatorname{sub}})$; i.e. $\omega(s+t)\leq\omega(s)+\omega(t)$ for all $s,t\geq 0$.
  • $(iv)$
    $\omega$ satisfies $(\omega_{\operatorname{nq}})$; i.e. $\int_{1}^{\infty}\frac{\omega(t)}{t^{2}}dt<+\infty$.
  • $(v)$
    $\omega$ satisfies $(\omega_{3,w})$; i.e. $\log(t)=O(\omega(t))$ as $t\rightarrow+\infty$.
Closely related to BB-weights are weights in the sense by H.-J. Petzsche and D. Vogt in [23]:
Definition 1.3.
According to [23] the function $\omega:[0,+\infty)\rightarrow[0,+\infty)$ is called a weight in the sense of Petzsche-Vogt, and write PV-weight for short, if
Note: The differences between BB- and PV-weights are subtle.
On the other hand in the Braun-Meise-Taylor setting the following set of functions is standard:
Definition 1.4.
    $\omega$ is a weight function in the sense of Definition 1.1,
    $\omega$ has $(\omega_{1})$; i.e. $\omega(2t)=O(\omega(t))$ as $t\rightarrow+\infty$,
    $\omega$ has $(\omega_{3})$; i.e. $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$,
    $\omega$ has $(\omega_{4})$; i.e. $\varphi_{\omega}:t\mapsto\omega(e^{t})$ is a convex function (on $\mathbb{R}$).
@@
Summarizing, when comparing the basic requirements on the weights it follows that subadditivity for $\omega$ in [1], [3] and [23] (Beurling-Björck setting) is replaced resp. relaxed in the Braun-Meise-Taylor setting by $\omega(2t)=O(\omega(t))$ which is frequently abbreviated by $(\alpha)$ in the literature. A further achievement is that the definition of weighted spaces within the Braun-Meise-Taylor setting admits the possibility to introduce and investigate the notion of quasianalyticity as well analogously to the weight sequence setting (“Denjoy-Carleman-Theorem”) whereas within the Beurling-Björck setting purely (non-trivial) test function spaces are studied; see Section 6.3 for precise definitions.
@@
On the other hand, for establishing the equivalent description in [7, Lemma 3.3, Prop. 3.4] crucially it has been used that $\varphi_{\omega}$ is convex; see $(\omega_{4})$ and this condition is frequently abbreviated by $(\delta)$ in the literature. $(\omega_{4})$ has become a standard assumption on $\omega$ also in related weighted settings since then; e.g. when considering globally defined functions of Gelfand-Shilov-type. This convexity condition is unavoidable for obtaining the coincidence of both settings since the proof of the Paley-Wiener-type theorem involves the double conjugate $\varphi^{**}_{\omega}$ and the identity $\varphi^{**}_{\omega}=\varphi_{\omega}$ which holds if and only if $\varphi_{\omega}$ is convex. In general, one only has $\varphi^{**}_{\omega}\leq\varphi_{\omega}$ and for this note that $\varphi^{**}_{\omega}$ and $\varphi^{*}_{\omega}$ both are automatically convex; see the comments in [28, Sect. 3.1] and more details are given in Section 2.9.
@@
In [28] and [24] we have shown that ultradifferentiable classes in the Braun-Meise-Taylor setting can equivalently be described by using the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}=(W^{(\ell)}_{j})_{j\in\mathbb{N}}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j))$; we refer to the main statement [24, Thm. 5.14] and for this alternative description $(\omega_{1})$ is crucial. It is immediate by definition that properties for $\omega$ can be transferred to $\mathcal{M}_{\omega}$ but, however, for the converse and hence when establishing a characterization of growth conditions, then again $(\omega_{4})$ is becoming relevant in order to treat the double conjugate. Similarly, this comment applies when properties for the classes $\mathcal{E}_{\{\omega\}}$, $\mathcal{E}_{(\omega)}$ are transferred to $\omega$ (resp. $\varphi_{\omega}$); see e.g. the proof of [24, Cor. 5.17]. Another crucial property for which $(\omega_{4})$ is unavoidable is the equivalence between $\omega$ and each associated weight function $\omega_{\mathbf{W}^{(\ell)}}$; see [24, Lemma 5.7], [28, Theorem 4.0.3, Lemma 5.1.3], and also [13, Lemma 2.5]. On the other hand, for the sake of completeness we mention that basically $\mathcal{M}_{\omega}$ can be introduced for more general $\omega$. Indeed, it suffices to assume that $\omega$ is non-decreasing and $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$ (i.e. $(\omega_{3})$) in order to ensure well-definedness of $\mathcal{M}_{\omega}$ (“matrix admissible weights”); see Section 2.9 for more details.
The main aim of this article is to go one step further, to define new weighted $L^{p}$-type spaces both of Roumieu- and Beurling-type in terms of abstractly given weight function matrices $\mathcal{W}:=\{\omega^{\ell}:\ell>0\}$ with matrix parameter $\ell>0$ (see Definition 3.1) and to answer the following question; see Sections 3, 4, 5 and 6:
@@
    Characterize the inclusion relations of weighted $L^{p}$-type spaces, and hence the equality (as l.c.v.s.), in terms of the defining matrices.
We answer $(Q1)$ in the main statements Theorem 4.1 for the Beurling-type and Theorem 4.2 for the Roumieu-type and work in a very general setting. First, we deal with abstractly given matrices $\mathcal{W}$ and, second, in Sections 3, 4 we are concerned with weighted subspaces of $L^{p}(\mathbb{R}^{d})$, $1\leq p\leq\infty$, subject to $\mathcal{W}$. Indeed, the weights are given by $x\mapsto\exp(\omega^{\ell}(x))$ and so involving the exponential function. Section 5 summarizes the situation when we focus on two special but natural weight function matrices for which the crucial information can be expressed in terms of a single weight function $\omega$: matrices of “exponential-type” and of “dilatation-type”. The special structure of these matrices allows for a more compact characterization. In Section 6 we apply the information to related and analogous situations: First, in Section 6.1 we consider the setting when weighting a fixed translation invariant subspace $\mathcal{A}^{p}\subseteq L^{p}(\mathbb{R}^{d})$. In Section 6.2 we focus in detail on the case when weighting the “Fourier image” of a given (fixed) modulation invariant function space $\mathcal{A}$; i.e. not the growth of $f$ directly is controlled but of its Fourier transform $\widehat{f}$. Of course, for this case it is crucial to ensure that the Fourier transform is well-defined on $\mathcal{A}$ and its image is non-trivial. Finally, in each instance we also study the anisotropic setting; i.e. when the weight functions are not radially extended to $\mathbb{R}^{d}$.
Summarizing, this general framework allows to treat in particular the known Beurling-Björck setting (for both types) which is corresponding to the special weight matrix $\omega^{\ell}:=\ell\omega$, $\ell>0$, where $\omega$ is given and fixed (“exponential type”), see Section 5.1, and when weighting the Fourier image of fixed subspaces (of functions having compact support) in $L^{1}(\mathbb{R}^{d})$. We refer to Section 6.3 for more details. Recall that $(Q1)$ for the Braun-Meise-Taylor setting has been solved in [24, Lemma 5.16 & Cor. 5.17] and the crucial growth relations are expressed by an O-growth resp. $o$-growth restriction and thus equivalence of weight functions characterizes the equality (as l.c.v.s.); see Section 2.3 for the relevant definitions. As expected, for the Beurling-Björck setting the same growth relations are becoming relevant.
In addition, we provide a detailed study concerning the coincidence of the new weighted spaces with the common (non-weighted) $L^{p}$-spaces (see Section 3.2), of their (non-)triviality (see Section 3.3) and investigate their translation invariance in Section 3.4.
By taking into account the information from $(Q1)$ and the fact that $(\omega_{4})$ is in general not preserved under equivalence of weight functions, we can see the second main question:
@@
    Illustrate the difference between the Beurling-Björck and the Braun-Meise-Taylor setting in view of the convexity for $\varphi_{\omega}$; more precisely we ask: How restrictive is $(\omega_{4})$? Does for any weight function (in the sense of Beurling-Björck) an equivalent function exist which satisfies $(\omega_{4})$?
Without $(\omega_{4})$ the main Paley-Wiener-type result from [7] and hence the equality between both settings is becoming unclear. Therefore, when treating the pure Beurling-Björck framework in view of $(Q1)$ the specific role resp. the loss of $(\omega_{4})$ has to be investigated. This question has been motivated by the explanations above and it is based on the comments in [28, Sect. 3.1]. But also in the recent article [20] in the introduction this convexity condition, its possible failure and its role in proofs within the Beurling-Björck are discussed. ($(\omega_{4})$ is also denoted by $(\delta)$ in [20].)
@@
The last Section 7 is dedicated to $(Q2)$ and to the construction of a (counter-)example showing that there exist weights in the sense of Beurling-Björck (and of Petzsche-Vogt) such that none equivalent (weight) function satisfies $(\omega_{4})$. Indeed, we even verify that there exist “many” of such weight functions; i.e. uncountable infinitely many different equivalence classes. The geometric construction also involves recent knowledge about the growth index $\gamma(\omega)$ for weight functions $\omega$ introduced in [12]; see Section 2.7. An immediate consequence of this example is to see the limitation of the weight sequence setting in this approach; i.e. when considering associated weight functions $\omega_{\mathbf{M}}$ with $\mathbf{M}$ being a sequence satisfying basic growth and regularity properties. This is due to the fact that $\omega_{\mathbf{M}}$ satisfies automatically $(\omega_{4})$; see Remark 7.6 for more details. On the other hand $\omega_{\mathbf{M}}$ appears frequently in the Braun-Meise-Taylor setting as a (counter-)example and in the comparison results from [6]; this again illustrates the difference between both settings. Due to the failure of $(\omega_{4})$ this (counter-)example also implies the fact that the information concerning the characterization of inclusion relations applies to spaces for which the corresponding BMT-setting results [24, Lemma 5.16 & Cor. 5.17] cannot be used. Finally, we emphasize that the constructed $\omega$ in Section 7 is matrix admissible.
@@
In the first Section 2 we list, summarize and compare in detail the numerous growth conditions and relations on weight functions in the different frameworks. One can expect that this content is also interesting for other different weighted settings since we exclusively deal with growth and regularity properties for functions. The same comment applies to the construction provided in Section 7.
Acknowledgements. The author of this article thanks the two anonymous referees for their efforts, the careful reading and the valuable suggestions which have improved the presentation of the results.
@@ 2.2. Relevant growth properties
The next condition is crucial for characterizing desired stability properties for Braun-Meise-Taylor classes $\mathcal{E}_{\{\omega\}}$ and $\mathcal{E}_{(\omega)}$ (of both types), see [24], [25] and the citations therein:
@@
In [28, $(3.8.4)$] this requirement has been denoted by $(\omega_{1^{{}^{\prime}}})$ but frequently in the literature it is abbreviated by $(\alpha_{0})$. Next, let us introduce the following two requirements appearing in the Beurling-Björck framework (see [3]):
Finally, recall that the Legendre-Fenchel-Young conjugate (of $\varphi_{\omega}$) is defined as follows (see (1.1)):
$\varphi^{*}_{\omega}(x):=\sup_{y\in\mathbb{R}}\{xy-\varphi_{\omega}(y)\},\;x\in\mathbb{R}.$
This transform is crucial when defining weighted classes in the sense of Braun-Meise-Taylor and it is also basic for introducing the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}=(W^{(\ell)}_{j})_{j\in\mathbb{N}}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp\left(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j)\right)$. We investigate this transform in more detail in Section 2.9.